Tohban Report 2014-07-23

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 16 Jul 2014
End Date: 23 Jul 2014
Tohban: Hazel Bain (Vice Tohban: Iain Hannah)
Tohban email:
Next Tohban: Mitsuo Oka
List all reports

Solar Activity

Solar activity this week was extremely quiet, with only 10 B class flares. The disk was almost spotless at points this week. The current GOES base level remains around the B1 level. NOAA 12121 produced two mid-B-level flares today but has otherwise shown no signs of development. Low level solar activity is expected for the next couple of days.

How many GOES flares occurred?

 Flares above B, C, M, X class were     10     0     0     0

RHESSI flare list contains 0 over the time range 16-Jul-14 23-Jul-14

Spacecraft Operations

The RHESSI cooldown from annealing continued this week with no apparent anomalies.

17-Jul-2014 19:20 UTC: The cryocooler power was increased slightly to 21 W in order to stay above 20 W. The cold plate was slowly cooling at a rate of about -0.5 degrees per orbit and the cold tip temp. was changing at a rate of about -0.6 degrees per orbit.

Cryo Power: 21.0 W avg accel: 15.7 mG

Cold Plate 1: 224.8 K Cold Plate 2: 223.1 K

Cold Tip 1: 187.7 K Cold Tip 2: 185.9 K

18-Jul-2014: The cold plate cooling rate was about -0.4 degrees per orbit and the cold tip temp was changing at a rate of about -0.2 degrees per orbit.

Cryo Power: 21.0 W avg accel: 14.6 mG

Cold Plate 1: 218.0 K Cold Plate 2: 216.2 K

Cold Tip 1: 183.3 K Cold Tip 2: 181.4 K

19-Jul-2014: The cold tip reached 180 K so the cryocooler power was increased to 22 W.

14-200-20:27:20 /icryomain power=68 14-200-20:28:28 /icryomain power=69

Cold Plate 1: 212 K Cold Plate 2: 210 K

Cold Tip 1: 178 K Cold Tip 2: 176 K Shroud: 216 K

Cryo Power: 22.4 W Vibrations: 13 mG

20-Jul-2014: The cold tips are going down at about 0.2 K per orbit.

Cold Plate 1: 206.8 K Cold Plate 2: 205.1 K

Cold Tip 1: 173.6 K Cold Tip 2: 171.9 K Shroud: 212.2 K

Cryo Power: 22.2 W Vibrations: 13.5 mG

21-Jul-2014: No cryocooler adjustments

Cold Plate 1: 201.9 K Cold Plate 2: 200.3 K

Cold Tip 1: 170.7 K Cold Tip 2: 168.9 K Shroud: 209.3 K

Cryo Power: 22.4 W Vibrations: 14.3 mG


The SSR is reaching a maximum of around 57%, and was getting down to around 18% at the end of the pass set yesterday. This is due to the increased the rate of SAS data during the anneal. (Since we are not taking detector data, it's a good time to record high-cadence SAS data.) For the time being this is not a concern but we should make sure we decrease the fill before we turn the detectors back on.

Personal tools